OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = Cube root of quantity x minus four.

OpenStudy (anonymous):

\(\Large f(x) = x^3 + 4\) \[\Large g(x) = \sqrt[3]{x-4}\] Let's do f(g(x)) and see if we get x. f(g(x)) means that we're taking g(x) and plugging it into f wherever there was an x.